\begin{thebibliography}{10}

\bibitem{barrios2014widder}
Bego{\~n}a Barrios, Ireneo Peral, Fernando Soria, and Enrico Valdinoci.
\newblock A {Widder's} type theorem for the heat equation with nonlocal
  diffusion.
\newblock {\em Archive for Rational Mechanics \& Analysis}, 213(2), 2014.

\bibitem{bhattacharyya2021inverse}
Sombuddha Bhattacharyya, Tuhin Ghosh, and Gunther Uhlmann.
\newblock Inverse problems for the fractional-laplacian with lower order
  non-local perturbations.
\newblock {\em Transactions of the American Mathematical Society},
  374:3053--3075, 2021.

\bibitem{cekic2020calderon}
Mihajlo Ceki{\'c}, Yi-Hsuan Lin, and Angkana R{\"u}land.
\newblock The {Calder\'on} problem for the fractional {Schr\"odinger} equation
  with drift.
\newblock {\em Calculus of Variations and Partial Differential Equations},
  59(91), 2020.

\bibitem{covi2020inverse}
Giovanni Covi.
\newblock An inverse problem for the fractional {Schr\"odinger} equation in a
  magnetic field.
\newblock {\em Inverse Problems}, 36(4):045004, 2020.

\bibitem{covi2020unique}
Giovanni Covi, Keijo M{\"o}nkk{\"o}nen, and Jesse Railo.
\newblock Unique continuation property and {Poincar\'e} inequality for higher
  order fractional laplacians with applications in inverse problems.
\newblock {\em Inverse Problems \& Imaging}, 15(4):641--681, 2021.

\bibitem{covi2020higher}
Giovanni Covi, Keijo M{\"o}nkk{\"o}nen, Jesse Railo, and Gunther Uhlmann.
\newblock The higher order fractional {Calder\'on} problem for linear local
  operators: uniqueness.
\newblock {\em arXiv preprint arXiv:2008.10227}, 2020.

\bibitem{dipierro2019local}
Serena Dipierro, Ovidiu Savin, and Enrico Valdinoci.
\newblock Local approximation of arbitrary functions by solutions of nonlocal
  equations.
\newblock {\em The Journal of Geometric Analysis}, 29(2):1428--1455, 2019.

\bibitem{felsinger2013local}
Matthieu Felsinger and Moritz Kassmann.
\newblock Local regularity for parabolic nonlocal operators.
\newblock {\em Communications in Partial Differential Equations},
  38(9):1539--1573, 2013.

\bibitem{fernandez2016boundary}
Xavier Fern{\'a}ndez-Real and Xavier Ros-Oton.
\newblock Boundary regularity for the fractional heat equation.
\newblock {\em Revista de la Real Academia de Ciencias Exactas, Fisicas y
  Naturales. Serie A. Matematicas}, 110(1):49--64, 2016.

\bibitem{ghosh2017calderon}
Tuhin Ghosh, Yi-Hsuan Lin, and Jingni Xiao.
\newblock The {Calder\'on} problem for variable coefficients nonlocal elliptic
  operators.
\newblock {\em Communications in Partial Differential Equations},
  42(12):1923--1961, 2017.

\bibitem{ghosh2020uniqueness}
Tuhin Ghosh, Angkana R{\"u}land, Mikko Salo, and Gunther Uhlmann.
\newblock Uniqueness and reconstruction for the fractional {Calder\'on} problem
  with a single measurement.
\newblock {\em Journal of Functional Analysis}, page 108505, 2020.

\bibitem{ghosh2020calderon}
Tuhin Ghosh, Mikko Salo, and Gunther Uhlmann.
\newblock The {Calder\'on} problem for the fractional {Schr\"odinger} equation.
\newblock {\em Analysis \& PDE}, 13(2):455--475, 2020.

\bibitem{greco2016existence}
Antonio Greco and Antonio Iannizzotto.
\newblock Existence and convexity of solutions of the fractional heat equation.
\newblock {\em Communications on Pure \& Applied Analysis}, 16(6):2201--2226,
  2017.

\bibitem{kian2020recovery}
Yavar Kian and Gunther Uhlmann.
\newblock Recovery of nonlinear terms for reaction diffusion equations from
  boundary measurements.
\newblock {\em arXiv preprint arXiv:2011.06039}, 2020.

\bibitem{krupchyk2020remark}
Katya Krupchyk and Gunther Uhlmann.
\newblock A remark on partial data inverse problems for semilinear elliptic
  equations.
\newblock {\em Proceedings of the American Mathematical Society},
  148(2):681--685, 2020.

\bibitem{kurylev2018inverse}
Yaroslav Kurylev, Matti Lassas, and Gunther Uhlmann.
\newblock Inverse problems for lorentzian manifolds and non-linear hyperbolic
  equations.
\newblock {\em Inventiones mathematicae}, 212(3):781--857, 2018.

\bibitem{lai2020inverse}
Ru-Yu Lai and Yi-Hsuan Lin.
\newblock Inverse problems for fractional semilinear elliptic equations.
\newblock {\em arXiv preprint arXiv:2004.00549}, 2020.

\bibitem{lai2020calderon}
Ru-Yu Lai, Yi-Hsuan Lin, and Angkana R{\"u}land.
\newblock The {Calder\'on} problem for a space-time fractional parabolic
  equation.
\newblock {\em SIAM Journal on Mathematical Analysis}, 52(3):2655--2688, 2020.

\bibitem{lai2021inverse}
Ru-Yu Lai and Ting Zhou.
\newblock An inverse problem for non-linear fractional magnetic schrodinger
  equation.
\newblock {\em arXiv preprint arXiv:2103.08180}, 2021.

\bibitem{lassas2021inverse}
Matti Lassas, Tony Liimatainen, Yi-Hsuan Lin, and Mikko Salo.
\newblock Inverse problems for elliptic equations with power type
  nonlinearities.
\newblock {\em Journal de math{\'e}matiques pures et appliqu{\'e}es},
  145:44--82, 2021.

\bibitem{li2021determining}
Li~Li.
\newblock Determining the magnetic potential in the fractional magnetic
  calder{\'o}n problem.
\newblock {\em Communications in Partial Differential Equations},
  46(6):1017--1026, 2021.

\bibitem{li2021fractional}
Li~Li.
\newblock A fractional parabolic inverse problem involving a time-dependent
  magnetic potential.
\newblock {\em SIAM Journal on Mathematical Analysis}, 53(1):435--452, 2021.

\bibitem{li2021inverse}
Li~Li.
\newblock On an inverse problem for a fractional semilinear elliptic equation
  involving a magnetic potential.
\newblock {\em Journal of Differential Equations}, 296:170--185, 2021.

\bibitem{liimatainen2020inverse}
Tony Liimatainen, Yi-Hsuan Lin, Mikko Salo, and Teemu Tyni.
\newblock Inverse problems for elliptic equations with fractional power type
  nonlinearities.
\newblock {\em arXiv preprint arXiv:2012.04944}, 2020.

\bibitem{miao2008well}
Changxing Miao, Baoquan Yuan, and Bo~Zhang.
\newblock Well-posedness of the {Cauchy} problem for the fractional power
  dissipative equations.
\newblock {\em Nonlinear Analysis: Theory, Methods \& Applications},
  68(3):461--484, 2008.

\bibitem{ruland2020fractional}
Angkana R{\"u}land and Mikko Salo.
\newblock The fractional {Calder\'on} problem: low regularity and stability.
\newblock {\em Nonlinear Analysis}, 193:111529, 2020.

\bibitem{ruland2020quantitative}
Angkana R{\"u}land and Mikko Salo.
\newblock Quantitative approximation properties for the fractional heat
  equation.
\newblock {\em Mathematical Control \& Related Fields}, 10(1):1--26, 2020.

\bibitem{silvestre2012holder}
Luis Silvestre.
\newblock {H\"older} estimates for advection fractional-diffusion equations.
\newblock {\em Annali della Scuola Normale Superiore di Pisa-Classe di
  Scienze}, 11(4):843--855, 2012.

\bibitem{sun1997inverse}
Ziqi Sun and Gunther Uhlmann.
\newblock Inverse problems in quasilinear anisotropic media.
\newblock {\em American journal of mathematics}, 119(4):771--797, 1997.

\bibitem{TaylorPDE3}
Michael Taylor.
\newblock {\em Partial Differential Equations III: Nonlinear Equations}.
\newblock Springer, 1996.

\bibitem{vazquez2014recent}
Juan~Luis V{\'a}zquez.
\newblock Recent progress in the theory of nonlinear diffusion with fractional
  laplacian operators.
\newblock {\em Discrete \& Continuous Dynamical Systems - S}, 7(4):857--885,
  2014.

\end{thebibliography}

\documentclass{article}
\usepackage[affil-it]{authblk}
\usepackage[utf8]{inputenc}
\usepackage{amsmath, amsfonts, amsthm, amssymb, latexsym}
\usepackage{geometry}
\usepackage{enumerate}
\usepackage{multirow}
\usepackage{rotating}
\usepackage{graphicx,color}
\usepackage{float}
\usepackage{mathrsfs}

\newtheorem{thm}{Theorem}[section]

\newtheorem{prop}[thm]{Proposition}

\newtheorem*{remark}{Remark}

\title{An inverse problem for a fractional diffusion equation with fractional power type nonlinearities}
\author{Li Li}
\affil{Institute for Pure and Applied Mathematics, University of California,\\
Los Angeles, CA 90095, USA}
\date{}

\begin{document}
\maketitle

\noindent \textbf{ABSTRACT.}\, 
We study the well-posedness of a semi-linear fractional diffusion equation and formulate an associated inverse problem.
We determine fractional power type nonlinearities from the exterior partial measurements of the Dirichlet-to-Neumann map. Our arguments are based on a first order linearization as well as the parabolic Runge approximation property.

\section{Introduction}
Non-local equations involving the fractional Laplacian $(-\Delta)^s$ have attracted much attention in past decades. These equations have been widely used for description of anomalous diffusion and random processes with jumps in many disciplines such as probability theory, physics and finance. 

Correspondingly, inverse problems associated with fractional operators involving $(-\Delta)^s$ have also been extensively studied so far. The study in this direction was initiated in \cite{ghosh2020calderon}. The authors considered 
the linear fractional elliptic problem
$$((-\Delta)^s+ q)u= 0\,\,\, \text{in}\,\,\Omega,\qquad
u|_{\Omega_e}= g$$
where $0< s< 1$. $\Omega\subset \mathbb{R}^n$ is a bounded domain with smooth boundary and $\Omega_e:= \mathbb{R}^n\setminus \bar\Omega$. They showed that the electric potential $q$ in $\Omega$ 
can be determined from 
the partial knowledge of the Dirichlet-to-Neumann map 
$$\Lambda_q: g\to (-\Delta)^su|_{\Omega_e}.$$
See \cite{bhattacharyya2021inverse, cekic2020calderon, covi2020inverse,  covi2020higher, ghosh2017calderon, ghosh2020uniqueness, li2021determining, ruland2020fractional} for further results for linear fractional elliptic inverse problems. 

In recent years, inverse problems associated with linear fractional parabolic operators have been studied as well. See \cite{lai2020calderon} for
the study of an inverse problem associated with the fractional operator 
$(\partial_t- \Delta)^s+ q$. See \cite{li2021fractional} for
the study of an inverse problem associated with a fractional parabolic operator involving time-dependent magnetic and electric potentials.

In this paper, we will study a fractional parabolic operator involving fractional power type nonlinearities and an associated inverse problem. In fact, the combination of fractional diffusion and power type nonlinearities has given rise to interesting mathematical models which have a number of scientific applications. The well-known equation
$$\partial_t u+ (-\Delta)^s(|u|^{m-1}u)= 0,$$
for instance, describes anomalous diffusion through porous media. See \cite{vazquez2014recent} for more details.

Here we will focus on another semi-linear fractional parabolic problem
\begin{equation}\label{fracparafrac}
\left\{
\begin{aligned}
\partial_t u+ (-\Delta)^s u + a(x, t, u)&= 0
\quad \,\,\,\Omega\times (0, T)\\
u&= g\quad \,\,\,\Omega_e\times (0, T)\\
u&= 0\quad \,\,\,\mathbb{R}^n\times \{0\}
\end{aligned}
\right.
\end{equation}
where the nonlinearity satisfies
\begin{equation}\label{fracpower}
a(x, t, z)= \sum^m_{k=1}a_k(x, t)|z|^{b_k}z,
 \end{equation}
$0\leq a_k\in C(\bar{\Omega}\times [0, T])$ and the powers $0< b_1<\cdots< b_m$ are not necessarily integers.

We will show (\ref{fracparafrac}) is well-posed at least for regular and small enough $g$, which enables us to define the parabolic Dirichlet-to-Neumann map
\begin{equation}\label{DNmap}
\Lambda_{a}g:= (-\Delta)^s u|_{\Omega_e\times (0, T)}.
\end{equation}
Our goal is to determine the nonlinearity from the exterior partial measurements of $\Lambda_{a}$. 

We mention that inverse problems associated with nonlinear equations have been extensively studied as well. See \cite{sun1997inverse} for inverse problems associated with quasi-linear elliptic 
equations. See \cite{kian2020recovery} for inverse problems associated with parabolic equations involving general semi-linear terms. See \cite{kurylev2018inverse} for inverse problems associated with nonlinear hyperbolic equations. We also remark that the higher order linearization technique has been commonly applied to determine the full nonlinearity in dealing with inverse problems associated with power type nonlinear equations. See \cite{krupchyk2020remark, lassas2021inverse, liimatainen2020inverse} for this approach for elliptic problems. See \cite{lai2020inverse, lai2021inverse} for the higher order linearization approach for fractional elliptic problems.

In this paper, we will only use a first order linearization to determine all the coefficients $a_k$ based on the Runge approximation property of the linear fractional parabolic operator. This approach can be viewed as a parabolic analogue of the one in \cite{li2021inverse}.

The following theorem is the main result in this paper.
\begin{thm}
Let $W_1, W_2\subset \Omega_e$ be open. Let $a^{(1)}, a^{(2)}$ be nonlinearities of the form (\ref{fracpower}). Suppose
$$\Lambda_{a^{(1)}}g|_{W_2\times (0, T)}
= \Lambda_{a^{(2)}}g|_{W_2\times (0, T)}$$
for small $g\in C^\infty_c(W_1\times (0, T))$.
Then $a^{(1)}_k= a^{(2)}_k$ in $\Omega\times(0, T)$, $k= 1,\cdots, m$.
\end{thm}

The rest of this paper is organized in the following way. In Section 2, we summarize the background knowledge. We show the well-posedness of  (\ref{fracparafrac}) in Section 3. We prove the main theorem in Section 4.~\\

\noindent \textbf{Acknowledgement.} The author would like to thank Professor Gunther Uhlmann for suggesting the problem and for helpful discussions.

\section{Preliminaries}

Throughout this paper we use the following notations.
\begin{itemize}
\item Fix the space dimension $n\geq 2$ and 
the fractional power $0< s< 1$.

\item Fix the constant $T> 0$ and $t$ denotes the time variable.

\item $\Omega$ denotes a bounded domain with smooth boundary and
$\Omega_e:= \mathbb{R}^n\setminus\bar{\Omega}$.

\item Let $u$ be an $(n+1)$-variable function. Then 
$u(t)$ denotes the $n$-variable function $u(\cdot, t)$.

\item $c, C, C', C_1,\cdots$ denote positive constants (which may depend on some parameters). We write $C_I$ when we emphasize that the constant $C$ depends on the parameter $I$.

\item $\langle\cdot, \cdot\rangle$ denotes the distributional pairing so formally, $\langle f, g\rangle= \int fg$.
\end{itemize}

\subsection{Function spaces}

Throughout this paper we refer all function spaces to real-valued function spaces.

For $r\in \mathbb{R}$, we have Sobolev spaces
$$H^r(\mathbb{R}^n):= \{f\in \mathcal{S}'(\mathbb{R}^n):
\int_{\mathbb{R}^n} (1+\vert \xi\vert^2)^r\vert \mathcal{F}f(\xi)\vert^2d\xi<\infty\}$$
where $\mathcal{F}$ is the Fourier transform and $\mathcal{S}'(\mathbb{R}^n)$ is the space of temperate distributions.

Let $U\subset \mathbb{R}^n$ be an open. 
We use $C^s(U)= C^{0, s}(U)$ to denote 
the H\"older space equipped with the standard norm
$$||f||_{C^s(U)}:= ||f||_{L^\infty(U)}+ 
\sup_{x\neq y, x,y\in U}\frac{|f(x)- f(y)|}{|x-y|^s}.$$
For $k\in \mathbb{N}$, the norm
$||\cdot||_{C^k(\mathbb{R}^n)}$ is defined by
$$||f||_{C^k(\mathbb{R}^n)}= 
\sum_{|\alpha|\leq k}||\partial^\alpha f||_{L^\infty(\mathbb{R}^n)}.$$

Let $X$ be a Banach space. We use $C([0, T]; X)$
to denote the space consisting of the corresponding Banach space-valued  continuous functions on $[0, T]$. 
$L^p(0, T; X)$ denotes the space consisting of the corresponding Banach space-valued $L^p$ functions, equipped with the standard norm
$$||u||_{L^p(0, T; X)}:= (\int^T_0||u(t)||^p_X\,dt)^{1/p}.$$

\subsection{The fractional Laplacian and the associated semi-group estimates}
For regular functions defined in $\mathbb{R}^n$, the non-local operator $(-\Delta)^s$ is given 
by the pointwise definition 
$$(-\Delta)^su(x):= 
c_{n,s}\lim_{\epsilon\to 0^+}
\int_{\mathbb{R}^n\setminus B_\epsilon(x)}\frac{u(x)-u(y)}{|x-y|^{n+2s}}\,dy$$
as well as the the equivalent bilinear form definition
$$\langle (-\Delta)^su, v\rangle:=  c'_{n,s}\int_{\mathbb{R}^n}\int_{\mathbb{R}^n}\frac
{(u(x)-u(y))(v(x)-v(y))}{|x-y|^{n+2s}}\,dxdy.$$

Let $\{S_\Omega(t)\}_{t\geq 0}$ be the semi-group on $L^2(\Omega)$ associated with the problem
\begin{equation}\label{bddsemigroup}
\left\{
\begin{aligned}
\partial_t u+ (-\Delta)^s u &= 0\quad \,\,\,\Omega\times \mathbb{R}_+\\
u&= 0\quad \,\,\,\Omega_e\times \mathbb{R}_+\\
u&= f\quad \,\,\,\mathbb{R}^n\times \{0\}.
\end{aligned}
\right.
\end{equation}
It has been proved in \cite{fernandez2016boundary} that for $f\in L^2(\Omega)$ and $t> 0$, $S_\Omega(t)f\in C^s(\mathbb{R}^n)$ and it is smooth in $t$. See \cite{felsinger2013local, silvestre2012holder} for more H\"older regularity results for fractional parabolic equations in bounded domains. Let $\{S(t)\}_{t\geq 0}$ be the semi-group associated with the problem
\begin{equation}\label{semigroup}
\left\{
\begin{aligned}
\partial_t u+ (-\Delta)^s u &= 0\quad \,\,\,\mathbb{R}^n\times \mathbb{R}_+\\
u&= f\quad \,\,\,\mathbb{R}^n\times \{0\}.
\end{aligned}
\right.
\end{equation}
By taking the Fourier transform, we have
$$S(t)f= K(\cdot, t)* f,\qquad K(x, t)= \int_{\mathbb{R}^n} e^{ix\cdot\xi- t|\xi|^{2s}}\,d\xi.$$
It has been shown that the associated kernel
$K(x, t)$ is smooth and positive in $\mathbb{R}^n\times \mathbb{R}_+$. See \cite{greco2016existence, miao2008well} for more details. The following estimate was given in Lemma 3.1 in \cite{miao2008well}.
\begin{prop}
Suppose $2\leq r\leq p\leq\infty$ and $f\in L^r(\mathbb{R}^n)$.
Then
$$||S(t)f||_{L^p(\mathbb{R}^n)}\leq Ct^{-\frac{n}{2s}(\frac{1}{r}- \frac{1}{p})}||f||_{L^r(\mathbb{R}^n)}.$$
\end{prop}

Now we identify $f\in L^r(\Omega)$ with its zero extension as a function in $L^r(\mathbb{R}^n)$ and
consider $\tilde{u}:= S(t)|f|- S_\Omega(t)|f|$. Note that $\tilde{u}$ solves the fractional parabolic equation in $\Omega\times \mathbb{R}_+$, $\tilde{u}\geq 0$ in $\Omega_e\times \mathbb{R}_+$ and $\tilde{u}(0)= 0$. By the fractional parabolic maximum principle (see for instance, Lemma 3.2 in \cite{barrios2014widder}), we have
$$|S_\Omega(t)f|\leq S_\Omega(t)|f|\leq S(t)|f|$$
in $\Omega\times \mathbb{R}_+$. Hence the following estimate immediately follows from the previous one.

\begin{prop}
Suppose $2\leq r\leq p\leq\infty$ and $f\in L^r(\Omega)$.
Then
$$||S_\Omega(t)f||_{L^p(\Omega)}\leq Ct^{-\frac{n}{2s}(\frac{1}{r}- \frac{1}{p})}||f||_{L^r(\Omega)}.$$
\end{prop}

\subsection{The Runge approximation property}
The following unique continuation property of $(-\Delta)^s$ was first established in Theorem 1.2 in \cite{ghosh2020calderon}. 
\begin{prop}
Let $0< s< 1$ and $u\in H^r(\mathbb{R}^n)$ for some $r\in \mathbb{R}$. Let $W\subset \mathbb{R}^n$ be open. If $$(-\Delta)^su= u= 0\quad\text{in}\,\,W,$$
then $u= 0$ in $\mathbb{R}^n$.
\end{prop}
This result was later extended for the fractional Laplacian when the fractional power belongs to $(-n/2, \infty)\setminus \mathbb{Z}$. See \cite{covi2020unique} for more details.

Based on this unique continuation property, the following parabolic Runge approximation property was established in Theorem 2 in \cite{ruland2020quantitative}.
\begin{prop}
Let $W\subset \Omega_e$ be open. Then the set
$$S:= \{u_g|_{\Omega\times (0, T)}: g\in C^\infty_c(W\times (0, T))\}$$
is dense in $L^2(\Omega\times (0, T))$. Here $u_g$ is the solution corresponding to the exterior data $g$ of the linear fractional parabolic problem
\begin{equation}\label{linearpara}
\left\{
\begin{aligned}
\partial_t u+ (-\Delta)^s u&= 0\quad \,\,\,\Omega\times (0, T)\\
u&= g\quad \,\,\,\Omega_e\times (0, T)\\
u&= 0\quad \,\,\,\mathbb{R}^n\times \{0\}.
\end{aligned}
\right.
\end{equation}
\end{prop}
This result was later extended for more general fractional parabolic operators (see Proposition 4.2 in \cite{li2021fractional}). See \cite{dipierro2019local} for $C^k$-type approximation results for fractional heat equations.

We remark that both properties are typical non-local phenomenons. They enable us to obtain strong results for inverse problems associated with fractional operators.

\section{The forward problem}
\subsection{Existence}
We say $(q, p, r)$ is admissible if $1< r\leq p< nr/(n-2s)$ and 
$$\frac{1}{q}= \frac{n}{2s}(\frac{1}{r}- \frac{1}{p}).$$
We also define
$$(Gf)(x, t):= \int^t_0 S_\Omega(t-\tau)f(x, \tau)\,d\tau.$$
We first prove the following two estimates, which are analogues of 
estimates established in Lemma 3.3 in \cite{miao2008well}.
\begin{prop}
Suppose $b>0$, $r> \frac{nb}{2s}$, $2(b+1)\leq  p< r(b+1)$
and $(q, p, r)$ is admissible. Then
$$||Gf||_{L^\infty(0, T; L^r(\Omega))}\leq CT^{1-\frac{nb}{2rs}}||f||_{L^{\frac{q}{b+1}}(0, T; L^\frac{p}{b+1}(\Omega))},$$
$$||Gf||_{L^q(0, T; L^p(\Omega))}\leq CT^{1-\frac{nb}{2rs}}||f||_{L^{\frac{q}{b+1}}(0, T; L^\frac{p}{b+1}(\Omega))}.$$
\end{prop}
\begin{proof}
First we note that $\frac{b+1}{q}< \frac{nb}{2rs}< 1$ since $(q, p, r)$ is admissible and $p< r(b+ 1)$.

To prove the first inequality, 
we apply Proposition 2.2 and H\"older inequality to obtain
$$||(Gf)(t)||_{L^r(\Omega)}\leq 
\int^t_0 ||S_\Omega(t-\tau)f(\tau, x)||_{L^r(\Omega)}\,d\tau$$
$$\leq C\int^t_0(t-\tau)^{-\frac{n}{2s}(\frac{b+1}{p}- \frac{1}{r})}
||f(\tau)||_{L^\frac{p}{b+1}(\Omega)}\,d\tau$$
$$\leq C(\int^t_0(t-\tau)^{-\frac{n}{2s}(\frac{b+1}{p}- \frac{1}{r})l}\,d\tau)^\frac{1}{l}||f||_{L^{\frac{q}{b+1}}(0, T; L^\frac{p}{b+1}(\Omega))}$$
$$= C't^{\frac{1}{l}-\frac{n}{2s}(\frac{b+1}{p}- \frac{1}{r})} ||f||_{L^{\frac{q}{b+1}}(0, T; L^\frac{p}{b+1}(\Omega))}
= C't^{1-\frac{nb}{2rs}}||f||_{L^{\frac{q}{b+1}}(0, T; L^\frac{p}{b+1}(\Omega))}$$
for each $t\in (0, T)$ where $l$ satisfies $\frac{1}{l}+ \frac{b+1}{q}= 1$. Hence we have
$$||(Gf)(t)||_{L^r(\Omega)}\leq C'T^{1-\frac{nb}{2rs}}||f||_{L^{\frac{q}{b+1}}(0, T; L^\frac{p}{b+1}(\Omega))}$$
for $t\in (0, T)$ since $1-\frac{nb}{2rs}> 0$ so the first inequality has been proved
.

To prove the second inequality, we apply Proposition 2.2 to obtain
$$||Gf||_{L^q(0, T; L^p(\Omega))}\leq 
||\int^t_0 ||S_\Omega(t-\tau)f(\tau, x)||_{L^p(\Omega)}\,d\tau||_{L^q}$$
$$\leq C||\int^t_0(t-\tau)^{-\frac{n}{2s}(\frac{b+1}{p}- \frac{1}{p})}
||f(\tau)||_{L^\frac{p}{b+1}(\Omega)}\,d\tau||_{L^q}$$

By Young's convolution inequality, for $t\in (0, T)$ we have
$$||\int^t_0(t-\tau)^{-\frac{nb}{2sp}}
||f(\tau)||_{L^\frac{p}{b+1}(\Omega)}\,d\tau||_{L^q}
\leq ||\tilde{f}* \tilde{g}||_{L^q}\leq ||\tilde{f}||_{L^\frac{q}{b+1}}||\tilde{g}||_{L^l}$$
$$=(\int^T_0 t^{-\frac{nb}{2ps}\cdot l}\,dt)^{\frac{1}{l}}||f||_{L^{\frac{q}{b+1}}(0, T; L^\frac{p}{b+1}(\Omega))}
= C'T^{1-\frac{nb}{2rs}}
||f||_{L^{\frac{q}{b+1}}(0, T; L^\frac{p}{b+1}(\Omega))}$$
where $l$ satisfies $1+ \frac{1}{q}= \frac{1+b}{q}+ \frac{1}{l}$, $\tilde{f}, \tilde{g}$
are defined to be zeros outside $(0, T)$ and for $t\in (0, T)$,
$$\tilde{f}(t):= ||f(t)||_{L^\frac{p}{b+1}(\Omega)},\quad 
\tilde{g}(t):= t^{-\frac{nb}{2ps}}.$$

\end{proof}

Now we consider the semi-linear fractional problem
\begin{equation}\label{nonhomopara}
\left\{
\begin{aligned}
\partial_t u+ (-\Delta)^s u + a(x, t, u)&= g\quad \,\,\,\Omega\times (0, T)\\
u&= 0\quad \,\,\,\Omega_e\times (0, T)\\
u&= 0\quad \,\,\,\mathbb{R}^n\times \{0\}
\end{aligned}
\right.
\end{equation}
where the nonlinearity has the form (\ref{fracpower}). 

\begin{remark}
Note that for fixed powers $0< b_1<\cdots< b_m$, we can choose a large $r$ s.t. $r> \max\{\frac{nb_m}{2s}, 2(b_m+ 1)\}$. For this chosen $r$, we can choose $p>r$ s.t.
$\frac{p}{r}< \min\{\frac{n}{n-2s}, b_1+1\}$. Then $p$ and $r$ satisfy all conditions in Proposition 3.1 for $b= b_k$ ($k= 1,\cdots,m$). 
\end{remark}

From now on we fix our choices for $p$ and $r$.

We define the nonlinear map 
$$(Fu)(x, t):= \int^t_0 S_\Omega(t-\tau)(g(x, \tau)- a(x, \tau, u(x, \tau)))\,d\tau$$
so the equation in (\ref{nonhomopara}) can be converted to the integral equation $u= Fu$ (see for instance, Section 15.1 in \cite{TaylorPDE3}).
\begin{prop}
For sufficiently small $g$, (\ref{nonhomopara}) has a solution in the space
$$X:= C([0, T]; L^r(\Omega))\cap L^q(0, T; L^p(\Omega)).$$
Here $X$ is equipped with the norm 
$$||\cdot||_X:= ||\cdot||_{L^q(0, T; L^p(\Omega))}+
||\cdot||_{L^\infty(0, T; L^r(\Omega))}.$$
\end{prop}
\begin{proof}
We will show $F$ is a contraction map on the complete metric space
$$X_M:=\{u\in X: ||u||_X\leq M\}.$$
Here the constant $M$ will be determined later, which depends on the norm of $g$.

Note that for $u\in X_M$, by Proposition 3.1 we have
$$||Fu||_X\leq CT^{1-\frac{nb_m}{2rs}}||g(x, t)- a(x,t,u(x,t))||_{L^{\frac{q}{b_m+1}}(0, T; L^\frac{p}{b_m+1}(\Omega))}.$$
Using H\"older inequalities we obtain
$$||a(x,t,u(x,t))||_{L^{\frac{q}{b_m+1}}(0, T; L^\frac{p}{b_m+1}(\Omega))}
\leq C'_{a, \Omega}||\sum^m_{k=1} ||u(t)||^{b_k+1}_{L^p(\Omega)}||_{L^\frac{q}{b_m+1}}$$
$$\leq C'_{a, \Omega, T}\sum^m_{k=1} ||u||^{b_k+1}_{L^q(0, T; L^p(\Omega))}
\leq C'_{a, \Omega, T}\sum^m_{k=1} M^{b_k+1}.$$
Hence we have
$$||Fu||_X\leq C''_{a, \Omega, T}(||g||_{L^{\frac{q}{b_m+1}}(0, T; L^\frac{p}{b_m+1}(\Omega))}+ \sum^m_{k=1} M^{b_k+1}).$$

Also note that
$$||z_2|^{b_k}z_2- |z_1|^{b_k}z_1|\leq C_k|z_2- z_1|(|z_1|^{b_k}+ |z_2|^{b_k})$$
so for $u_1, u_2\in X_M$, by Proposition 3.1 and H\"older inequalities
we get that 
$$||Fu_1- Fu_2||_X\leq CT^{1-\frac{nb_m}{2rs}}||a(x,t,u_2(x,t))- a(x,t,u_1(x,t))||_{L^{\frac{q}{b_m+1}}(0, T; L^\frac{p}{b_m+1}(\Omega))}$$
$$\leq C_{a, \Omega, T}||u_1- u_2||_{L^q(0, T; L^p(\Omega))}
\sum^m_{k=1}(||u_1||^{b_k}_{L^q(0, T; L^p(\Omega))}
+ ||u_2||^{b_k}_{L^q(0, T; L^p(\Omega))})$$
$$\leq 2C_{a, \Omega, T}||u_1- u_2||_X\sum^m_{k=1} M^{b_k}.$$

Now for sufficiently small $g$, we choose $M$ to be $||g||^{1/2}_{L^{\frac{q}{b+1}}(0, T; L^\frac{p}{b+1}(\Omega))}$. 

Then
$||Fu||_X\leq M$ for $u\in X_M$ and $||Fu_1- Fu_2||_X\leq \frac{1}{2}||u_1- u_2||_X$ for $u_1, u_2\in X_M$. We can estimate $(Fu)(x, t+h)- (Fu)(x, t)$ in similar ways to show the continuity of $Fu$ in $t$. Hence $F$ has a fix point
in $X_M$ by the fixed-point theorem.
\end{proof}

Note that $||(-\Delta)^s f||_{L^\infty(\mathbb{R}^n)}\leq C||f||_{C^2(\mathbb{R}^n)}$ for $f\in C^\infty_c(\mathbb{R}^n)$
(see for instance, Lemma 3.3 in \cite{li2021inverse}) so by Proposition 3.2 the problem
\begin{equation}
\left\{
\begin{aligned}
\partial_t w+ (-\Delta)^s w + a(x, t, w)&= -(-\Delta)^s g\quad \,\,\,\Omega\times (0, T)\\
w&= 0\qquad\qquad \,\,\,\,\,\,\,\Omega_e\times (0, T)\\
w&= 0\qquad\qquad \,\,\,\,\,\,\,\mathbb{R}^n\times \{0\}
\end{aligned}
\right.
\end{equation}
has a solution in $X$ for small $g\in C^\infty_c(\Omega_e\times (0, T))$. Then $u:= w+ g$ gives a solution of (\ref{fracparafrac}).

\subsection{$L^\infty$-estimate and uniqueness}
We first prove the following $L^\infty$-estimate, which will be used in the first order linearization later.

\begin{prop}
Suppose $u$ is a solution of the problem
\begin{equation}\label{fgpara}
\left\{
\begin{aligned}
\partial_t u+ (-\Delta)^s u + a(x, t, u)&= f
\quad \,\,\,\Omega\times (0, T)\\
u&= g\quad \,\,\,\Omega_e\times (0, T)\\
u&= 0\quad \,\,\,\mathbb{R}^n\times \{0\}.
\end{aligned}
\right.
\end{equation}
\end{prop}
Then we have
$$||u||_{L^\infty}\leq T||f||_{L^\infty(\Omega\times (0, T))}
+ ||g||_{L^\infty(\Omega_e\times (0, T))}.$$
\begin{proof}
We fix $\phi\in C^\infty_c(\mathbb{R}^n)$ s.t. $0\leq \phi\leq 1$
and $\phi= 1$ on $\bar\Omega \cup \bar W$. We define
$$\tilde{\phi}(x, t):= (||f||_{L^\infty(\Omega\times (0, T))}t
+ ||g||_{L^\infty(\Omega_e\times (0, T))})\phi(x).$$
Clearly $(-\Delta)^s\phi\geq 0$ in $\Omega$ from the pointwise definition 
of $(-\Delta)^s$ so $$\partial_t\tilde{\phi}+(-\Delta)^s\tilde{\phi}+ a(x, t, \tilde{\phi})\geq ||f||_{L^\infty(\Omega\times (0, T))}$$
in $\Omega\times (0, T)$. Now we consider $\tilde{u}:= \tilde{\phi}- u$.
Note that $\tilde{u}\geq 0$ in 
$\Omega_e\times (0, T)$, $\tilde{u}\geq 0$ at $t= 0$ and 
$$\partial_t\tilde{u}+(-\Delta)^s\tilde{u}+ a(x, t, \tilde{\phi})
- a(x, t, u)\geq 0$$
in $\Omega\times (0, T)$. Write $\tilde{u}= \tilde{u}^+ - \tilde{u}^-$
where $\tilde{u}^{\pm}= \max\{\pm \tilde{u}, 0\}$. Then $\tilde{u}^-= 0$ in $\Omega_e\times (0, T)$. 

Now we define $E(t):= ||\tilde{u}^-(t)||^2_{L^2(\Omega)}
\geq 0$. Then 
$E(0)= 0$ and
$$E'(t)= -2\langle \partial_t\tilde{u}(t), \tilde{u}^-(t)\rangle
\leq 2\langle (-\Delta)^s\tilde{u}(t), \tilde{u}^-(t)\rangle
+ 2\langle a(x, t, \tilde{\phi})- a(x, t, u), \tilde{u}^-(t)\rangle.$$
Note that
$$\langle (-\Delta)^s\tilde{u}(t), \tilde{u}^-(t)\rangle=
\iint\frac{(\tilde{u}(t)(x)- \tilde{u}(t)(y))(\tilde{u}^-(t)(x)- \tilde{u}^-(t)(y))}{|x-y|^{n+2s}}dxdy.$$
Since $(\tilde{u}^+(t)(x)- \tilde{u}^+(t)(y))(\tilde{u}^-(t)(x)- \tilde{u}^-(t)(y))\leq 0$, we have 
$$\langle (-\Delta)^s\tilde{u}(t), \tilde{u}^-(t)\rangle
\leq -\iint\frac{(\tilde{u}^-(t)(x)- \tilde{u}^-(t)(y))^2}{|x-y|^{n+2s}}dxdy\leq 0.$$
Also note that $a(x, t, \tilde{\phi})- a(x, t, u)$ have the same sign as $\tilde{u}$. Hence
we have $E'(t)\leq 0$ so the only possibility is $E(t)= 0$ and thus $\tilde{\phi}\geq u$ in $\Omega\times (0, T)$. 

Also note that $a(x, t, \tilde{\phi})+ a(x, t, u)$ have the same sign as $\tilde{\phi}+ u$ so similarly we can consider $\tilde{u}:= \tilde{\phi}+ u$ and show $\tilde{\phi}\geq -u$ in $\Omega\times (0, T)$. Hence we have
$|u|\leq \tilde{\phi}$ in $\Omega\times (0, T)$.
\end{proof}

\begin{remark}
In fact, we can use similar arguments to show the uniqueness of solutions of (\ref{fgpara}). 
Suppose $u_1, u_2$ are two solutions of (\ref{fgpara}). Let $\tilde{u}:= u_1- u_2$. Then
$\tilde{u}= 0$ in 
$\Omega_e\times (0, T)$, $\tilde{u}= 0$ at $t= 0$ and 
$$\partial_t\tilde{u}+(-\Delta)^s\tilde{u}+ a(x, t, u_1)
- a(x, t, u_2)= 0$$
in $\Omega\times (0, T)$. Since $a(x, t, u_1)
- a(x, t, u_2)$ has the same sign as $\tilde{u}$, we can show that 
$$E'(t)= -2\langle \partial_t\tilde{u}(t), \tilde{u}^-(t)\rangle
= 2\langle (-\Delta)^s\tilde{u}(t), \tilde{u}^-(t)\rangle
+ 2\langle a(x, t, u_1)- a(x, t, u_2), \tilde{u}^-(t)\rangle\leq 0$$
where
$E(t):= ||\tilde{u}^-(t)||^2_{L^2(\Omega)}$ as before to conclude that $u_1\geq u_2$.
 Similarly we consider $\tilde{u}= u_2- u_1$ to prove $u_2\geq u_1$. Hence $u_1= u_2$. 
\end{remark}

In particular, we conclude that (\ref{fracparafrac}) has a unique solution.

\section{The inverse problem}
The well-posedness result in the previous section ensures that the Dirichlet-to-Neumann map
$$\Lambda_{a}g:= (-\Delta)^s u|_{\Omega_e\times (0, T)}$$
is well-defined at least for regular and small $g$.

\begin{remark}
It has been shown that the knowledge of $\Lambda_{a}$ is equivalent to the knowledge of the non-local Neumann operator $\mathcal{N}_su$ (see for instance, \cite{ghosh2020calderon}), which is defined by
$$\mathcal{N}_su(t)(x):= 
c_{n,s}
\int_{\Omega}\frac{u(t)(x)-u(t)(y)}{|x-y|^{n+2s}}\,dy,\quad 
x\in \Omega_e, t\in (0, T).$$ 
Recall that for the inverse problem for the classical semi-linear parabolic operator (see for instance, \cite{kian2020recovery}), the 
Dirichlet-to-Neumann map was defined  by
$$\Lambda_{a}: g\to \partial_\nu u|_{\partial\Omega\times (0, T)}$$
where $\partial_\nu u$ is the classical Neumann derivative associated with the 
initial boundary value problem
\begin{equation*}
\left\{
\begin{aligned}
\partial_t u -\Delta u + a(x, t, u)&= 0\quad \,\,\,\Omega\times (0, T)\\
u&= g\,\,\,\,\, \partial\Omega\times (0, T)\\
u&= 0\quad \,\,\,\Omega\times \{0\}.
\end{aligned}
\right.
\end{equation*}
Hence our inverse problem can be viewed as a natural non-local analogue of the classical problem.
\end{remark}

We will apply the linearization scheme, which enables us to use the Runge approximation property for the linear fractional parabolic operator to deal with the inverse problem for the 
semi-linear fractional parabolic operator.
 
\subsection{Linearization}
For $g\in C^\infty_c(\Omega_e\times (0, T))$ and small $\lambda> 0$, we use 
$u_g$ to denote the solution of the linear problem (\ref{linearpara}) and we use $u_{\lambda, g}$ to denote the solution of the semi-linear problem
\begin{equation}\label{lambdafrac}
\left\{
\begin{aligned}
\partial_t u+ (-\Delta)^s u + a(x, t, u)&= 0\quad \,\,\,\Omega\times (0, T)\\
u&= \lambda g\,\,\,\,\, \Omega_e\times (0, T)\\
u&= 0\quad \,\,\,\mathbb{R}^n\times \{0\}.
\end{aligned}
\right.
\end{equation}

\begin{prop}
Let $v_{\lambda, g}:= u_g- \frac{u_{\lambda, g}}{\lambda}$. Then $\lim_{\lambda\to 0}v_{\lambda, g}= 0$
in $L^\infty(\Omega\times (0, T))$.
\end{prop}
\begin{proof}
Note that $v_{\lambda, g}= 0$ in $\Omega_e\times (0, T)$ and
$$\partial_t v_{\lambda, g}+ (-\Delta)^s v_{\lambda, g}= \frac{1}{\lambda}
a(x, t, u_{\lambda, g})$$
in $\Omega\times (0, T)$. By the Proposition 3.3, we have
$$||v_{\lambda, g}||_{L^\infty}\leq \frac{T}{\lambda}
||a(x, t, u_{\lambda, g})||_{L^\infty(\Omega\times (0, T))}$$
and $||u_{\lambda, g}||_{L^\infty}\leq 
\lambda||g||_{L^\infty(\Omega_e\times (0, T))}$ so
$$||v_{\lambda, g}||_{L^\infty}\leq 
T\sum^m_{k=1}\lambda^{b_k}||a_k(x,t)||_{L^\infty}||g||^{b_k+1}_{L^\infty(\Omega_e\times (0, T))},$$
which implies $||v_{\lambda, g}||_{L^\infty}\to 0$ as $\lambda\to 0$.
\end{proof}

\subsection{Proof of the main theorem}
Now we are ready to prove Theorem 1.1.
\begin{proof}
Since we have the assumption
$$\Lambda_{a^{(1)}}g|_{W_2\times (0, T)}
= \Lambda_{a^{(2)}}g|_{W_2\times (0, T)}$$
for small $g\in C^\infty_c(W_1\times (0, T))$ and 
$u^{(1)}_{\lambda, g}= u^{(2)}_{\lambda, g}= \lambda g$ in $\Omega_e\times (0, T)$,
for each $t$ we have
$$(-\Delta)^s(u^{(1)}_{\lambda, g}- u^{(2)}_{\lambda, g})(t)|_{W_2}= 0, \qquad (u^{(1)}_{\lambda, g}- u^{(2)}_{\lambda, g})(t))|_{\Omega_e}= 0$$
for $g\in C^\infty_c(W_1\times (0, T))$ and small $\lambda> 0$ so Proposition 2.3 implies 
$u^{(1)}_{\lambda, g}= u^{(2)}_{\lambda, g}=:  u_{\lambda, g}$ in $\mathbb{R}^n\times (0, T)$. Hence we have
$$a^{(1)}(x, t, u_{\lambda, g}(x,t))
= a^{(2)}(x, t, u_{\lambda, g}(x,t))$$
\begin{equation}\label{aRid}
(a^{(1)}_1(x, t)- a^{(2)}_1(x, t))|u_{\lambda, g}|^{b_1}u_{\lambda, g}
= R^{(2)}_1(x, t, u_{\lambda, g})
- R^{(1)}_1(x, t, u_{\lambda, g})
\end{equation}
in $\Omega\times (0, T)$ where
$$R^{(i)}_j(x, t, z):= \sum^m_{k=j+1}a^{(i)}_k(x, t)|z|^{b_k}z.$$

Now note that
$$|||a^{(1)}_1(x,t)- a^{(2)}_1(x,t)|^\frac{1}{b_1+1}||_{L^2(\Omega\times 
(0, T))}$$
$$\leq |||a^{(1)}_1(x, t)- a^{(2)}_1(x, t)|^\frac{1}{b_1+1}(1- \frac{u_{\lambda, g}}{\lambda})||_{L^2(\Omega\times (0, T)}$$
$$+ \frac{1}{\lambda}|||a^{(1)}_1(x,t)- a^{(2)}_1(x,t)|^\frac{1}{b_1+1}
u_{\lambda, g}||_{L^2(\Omega\times (0, T))}$$
$$\leq |||a^{(1)}_1(x, t)- a^{(2)}_1(x, t)|^\frac{1}{b_1+1}||_{L^\infty}||1- \frac{u_{\lambda, g}}{\lambda}||_{L^2(\Omega\times (0, T))}$$
\begin{equation}\label{2inftyest}
+ \frac{1}{\lambda}|||a^{(1)}_1(x,t)- a^{(2)}_1(x,t)|^\frac{1}{b_1+1}
u_{\lambda, g}||_{L^2(\Omega\times (0, T))}.
\end{equation}
For given $\delta> 0$, by Proposition 2.4 we can choose 
$g\in C^\infty_c(W_1\times (0, T))$ s.t. 
$$||1- u_g||_{L^2(\Omega\times (0, T))}\leq \delta$$ 
and for this chosen $g$, we have 
\begin{equation}\label{Rungelinear}
||1- \frac{u_{\lambda, g}}{\lambda}||_{L^2(\Omega\times (0, T))}\leq 2\delta
\end{equation}
for small $\lambda$ by Proposition 4.1. Since Proposition 3.3 implies that
$$||u_{\lambda, g}||_{L^\infty}\leq 
\lambda||g||_{L^\infty(\Omega_e\times (0, T))},$$
by (\ref{aRid}) we have
$$\frac{1}{\lambda}|||a^{(1)}_1(x,t)- a^{(2)}_1(x,t)|^\frac{1}{b_1+1}
u_{\lambda, g}||_{L^2(\Omega\times (0, T))}$$
$$\leq 
\frac{C}{\lambda}|||R^{(2)}_1(x, t, u_{\lambda, g})
- R^{(1)}_1(x, t, u_{\lambda, g})|^{\frac{1}{b_1+1}}||_{L^\infty(\Omega\times (0, T))}$$
$$\leq C'(\sum^m_{k=2}\lambda^{\frac{b_k-b_1}{b_1+1}}(||a^{(1)}_k||^
\frac{1}{b_1+1}_{L^\infty}+||a^{(2)}_k||^
\frac{1}{b_1+1}_{L^\infty})||g||^\frac{b_k+1}{b_1+1}_{L^\infty}).$$
This inequality implies
$$\frac{1}{\lambda}|||a^{(1)}_1(x,t)- a^{(2)}_1(x,t)|^\frac{1}{b_1+1}
u_{\lambda, g}||_{L^2(\Omega\times (0, T))}\to 0$$
as $\lambda\to 0$. Then by (\ref{2inftyest}) and (\ref{Rungelinear}) we obtain
$$|||a^{(1)}_1(x,t)- a^{(2)}_1(x,t)|^\frac{1}{b_1+1}||_{L^2(\Omega\times (0, T))}\leq 
2\delta |||a^{(1)}_1(x,t)- a^{(2)}_1(x,t)|^\frac{1}{b_1+1}||_{L^\infty(\Omega\times (0, T))}.$$
Now we conclude that $a^{(1)}_1= a^{(2)}_1$ since $\delta$ is arbitrary.

Iteratively, once we have shown $a^{(1)}_j= a^{(2)}_j$ ($1\leq j\leq m'-1$), we have 
$$(a^{(1)}_{m'}(x, t)- a^{(2)}_{m'}(x, t))|u_{\lambda, g}|^{b_{m'}}
u_{\lambda, g}
= R^{(2)}_{m'}(x, t, u_{\lambda, g})
- R^{(1)}_{m'}(x, t, u_{\lambda, g})$$
in $\Omega\times (0, T)$. Also note that
$$|||a^{(1)}_{m'}(x,t)- a^{(2)}_{m'}(x,t)|^\frac{1}{b_{m'}+1}||_{L^2(\Omega\times 
(0, T))}$$
$$\leq |||a^{(1)}_{m'}(x, t)- a^{(2)}_{m'}(x, t)|^\frac{1}{b_{m'}+1}||_{L^\infty}||1- \frac{u_{\lambda, g}}{\lambda}||_{L^2(\Omega\times (0, T))}$$
$$+ \frac{1}{\lambda}|||a^{(1)}_{m'}(x,t)- a^{(2)}_{m'}(x,t)|^\frac{1}{b_{m'}+1}
u_{\lambda, g}||_{L^2(\Omega\times (0, T))}.$$
For given $\delta> 0$, we can choose 
$g\in C^\infty_c(W_1\times (0, T))$ s.t. 
$$||1- \frac{u_{\lambda, g}}{\lambda}||_{L^2(\Omega\times (0, T))}\leq 2\delta$$
for small $\lambda$ and we also have
$$\frac{1}{\lambda}|||a^{(1)}_{m'}(x,t)- a^{(2)}_{m'}(x,t)|^\frac{1}{b_{m'}+1}
u_{\lambda, g}||_{L^2(\Omega\times (0, T))}$$
$$\leq 
\frac{C}{\lambda}|||R^{(2)}_{m'}(x, t, u_{\lambda, g})
- R^{(1)}_{m'}(x, t, u_{\lambda, g})|^{\frac{1}{b_{m'}+1}}||_{L^\infty(\Omega\times (0, T))}$$
$$\leq C'(\sum^m_{k={m'}+1}\lambda^{\frac{b_k-b_{m'}}{b_{m'}+1}}(||a^{(1)}_k||^
\frac{1}{b_{m'}+1}_{L^\infty}+||a^{(2)}_k||^
\frac{1}{b_{m'}+1}_{L^\infty})||g||^\frac{b_k+1}{b_{m'}+1}_{L^\infty}).$$
Now let $\lambda\to 0$. Then we get
$$|||a^{(1)}_{m'}(x,t)- a^{(2)}_{m'}(x,t)|^\frac{1}{b_{m'}+1}||_{L^2(\Omega\times (0, T))}\leq 
2\delta |||a^{(1)}_{m'}(x,t)- a^{(2)}_{m'}(x,t)|^\frac{1}{b_{m'}+1}||_{L^\infty(\Omega\times (0, T))}.$$
Now we conclude that $a^{(1)}_{m'}= a^{(2)}_{m'}$ since $\delta$ is arbitrary.
\end{proof}

\bibliographystyle{plain}
{\small\bibliography{Reference4}}
\end{document}